OpenStudy (anonymous):

A ball is thrown upward with an initial velocity of 48 ft/s from a height 864 ft. h = -16t +48t +864 After how many seconds will the ball reach the ground? (i.e. For what positive value of t is h=0?)

OpenStudy (anonymous):

are you sure it is not h = -16t^2 +48t +864 ?

OpenStudy (anonymous):

It is! Sorry, bad error!

OpenStudy (anonymous):

0 = -16t^2 +48t +864 solve for t

OpenStudy (anonymous):

you can divide out by 16

OpenStudy (anonymous):

so it'd be -t^2 + 3t + 54 ?

OpenStudy (anonymous):

-t^2 + 3t + 54=0 yes

OpenStudy (anonymous):

-t^2 + 3t + 54=0 you can even divide by -1 if you want, t^2-3t-54=0

OpenStudy (anonymous):

do you then factor it?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i can't think of any numbers that would work for that particular factoring :/

OpenStudy (anonymous):

two number that add up to -3, and multiply into -54?

OpenStudy (anonymous):

-9 , 6 -9+6=-3 -9*6=-54

OpenStudy (anonymous):

(x-9)(x+6)=0

OpenStudy (anonymous):

so the answer would be, the ball will reach the ground in 9 seconds?